OpenStudy (anonymous):

6.78j-5.2=4.33j+2.15 j =3 how is this?

OpenStudy (anonymous):

substitute 3 in every place for j:\[6.78j-5.2=4.33j+2.15~~\implies~~ 6.78(3)-5.2=4.33(3)+2.15\]

OpenStudy (asnaseer):

the way to solve this is to collect all terms involving j on the left-hand-side of the equals sign, and all terms that are just numbers on the right-hand-side

OpenStudy (asnaseer):

do that first and see what you get

OpenStudy (anonymous):

\[ 6.78(3)−5.2=4.33(3)+2.15~~\implies~~20.34-5.2=12.99+2.15\]\[~~~~~20.34-5.2=12.99+2.15~~\implies~~15.14=15.14\]

OpenStudy (asnaseer):

that is not how you solve this type of equation to find j.